An automobile company sells 5 trucks for every 9 sedans. If the company sells 15 trucks, how many sedans will they sell?
Answer by jim_thompson5910(35256)

(# of trucks)/(# of sedans) = (# of trucks)/(# of sedans)
5/9 = 15/x
5x = 9*15
5x = 135
5x/5 = 135/5
x = 27

Question: If the company sells 15 trucks, how many sedans will they sell?
Answer: 27 sedans

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------